if p,r,and s are three different prime numbers greater than 2, and n= p x r x s, how many positive factors, including 1 and n, does n have?
Answer by robertb(5830)

The power of p is 1, the power of r is 1, and the power of s is 1, so the total number of divisors, or factors, of n, is (1+1)(1+1)(1+1) = 2*2*2 = 8.